# TASR: Timestep-Aware Diffusion Model for Image Super-Resolution

### v1

@@ 4 Experiments
4.1 Dataset and Evaluation Metric
@@
Datasets. Following [43, 46, 16], we train our method on DIV2K [2], DIV8K [10], Flickr2K [35], OST [38], and the first 10,000 face images from FFHQ [13], and use the degradation pipeline of RealESR-GAN [40] to generate HR-LR image pairs for training. We evaluate the performance of our method on both synthetic and real-world datasets. The synthetic dataset is generated from the DIV2K validation set, where we use the same degradation pipeline in the training process to randomly crop 3K image patches to as LR images. For the real-world test datasets, we evaluate the DrealSR [42] and RealSR [42] datasets, where each image is center-cropped to obtain LR images of resolution. The resolution of each HR image in both the training and test sets is .
@@
Metrics. We perform a comprehensive and effective quantitative evaluation of ISR methods using a series of widely used reference and non-reference metrics. Among the reference-based metrics, PSNR and SSIM [41] (calculated on the Y channel in the YCbCr space) are fidelity metrics, while LPIPS [51] are quality assessment metrics. MANIQA [45], MUSIQ [45], and CLIPIQA [36] are non-reference image quality assessment (IQA) metrics.
4.2 Implementation Details
@@
We use the pre-trained SD-v2.1 [28] model as the base SD model. In the first stage of training, we fine-tune the ControlNet module for 20K iterations, and in the second stage, we fine-tune both the IRControlNet and Adapter for 100K iterations. During training, we use the AdamW [14] optimizer with a weight decay of 1e-2, a batch size of 32, and a learning rate of 1e-5. All experiments are conducted at a resolution of 512×512 on 8 NVIDIA A100 GPUs. During inference, we employ a classifier-free guidance strategy, generating higher-quality images through negative prompts without additional training. The guidance scale for classifier-free guidance is set to 4.5, and we use a spaced DDPM sampling schedule [21] with 20 timesteps.
4.3 Comparison with State-of-the-art Methods
@@
To validate the effectiveness of our method, we compared it with several state-of-the-art GAN-based and diffusion-based ISR methods, namely RealESRGAN [40], BSRGAN [49], SwinIR [15], SeeSR [43], PASD [46], ResShift [48], DiffBIR [16], SUPIR [47].
@@
Quantitative Comparisons. Tab. 1 presents the quantitative comparison on both synthetic and real-world test datasets. As observed, our method significantly outperforms other state-of-the-art (SOTA) methods on non-reference metrics such as MANIQA, MUSIQ, and CLIPIQA across all datasets, generating higher-quality HR images. For example, on the DRealSR dataset, TASR outperforms the second-best method, SeeSR by 7.8%, 0.4%, and 5.0% on MANIQA, MUSIQ, and CLIPIQA metrics, respectively. Furthermore, GAN-based methods surpass diffusion-based methods on reference metrics such as PSNR and SSIM. We attribute this phenomenon to the fact that diffusion-based methods leverage powerful generative priors to generate details that are more perceptually realistic to humans, but this comes at the expense of the fidelity to the LR images, as noted in previous works [43, 46].
@@ 4.4 Ablation Study
Model Architecture. We employ various architectures to validate the effectiveness of our proposed model structure for the adapter. Firstly, we removed all time-adaptive normalization layers from the adapter and directly utilized the U-Net features and the skip connection features to predict the control weight map . This modification is denoted as ‘w/o timestep’. As shown in Tab. 2, by incorporating timesteps into the adapter, our method achieves better performance in all the metrics compared to the variant without timesteps. In addition, we replace our adapter structure with a transformer-based architecture similar to ELLA [12], denoted as ‘Transformer’. The variant with the Transformer block has declined in the non-reference metrics.
@@
Training for Different Module. Firstly, we remove the timestep-aware adapter and optimize only the parameters of ControlNet during the training process, denoted as ‘ControlNet’. As shown in Tab. 3 and Fig. 5, when the limitation of the adapter on the optimization space is removed, the method is prone to reward hacking during training. Although there is a significant improvement in perceptual metrics such as CLIPIQA and MUSIQ, the generated images tend to align with the preferences of these metrics rather than human perception. This can result in images that are not sufficiently realistic and lack fidelity, such as the strange feathers and the human eye in the sky background in Fig. 5 (a). Similarly, we add the proposed adapter but optimize only the parameters of the adapter during the training process, denoted as ‘Adapter’. Our method shows a better trade-off between the reference and non-reference metrics. When only the Adapter is optimized, the parameter space available for optimization is constrained, resulting in a 6.3% and 2.0% decrease in the MANIQA, and MUSIQ metrics, respectively. As can be seen from Fig. 5, compared to our method, the model obtained by optimizing only the Adapter struggles to generate more refined feathers.
@@
Loss Function. We begin by evaluating the impact of the chosen loss functions on the results. As shown in Tab. 4, when only is added as the loss function (row 2), the no-reference image quality assessment metrics MANIQA improves by 8.3% compared to using only the denoising loss (row 1). However, reference metrics such as PSNR and SSIM decrease by 4.3% and 10.6% respectively, indicating lower fidelity in the generated HR images. Similarly, when only is added as the loss function (row 3), the reference metrics improve while the non-reference metrics correspondingly decline. In contrast, our proposed training strategy, applying and , allows us to enhance image perceptual quality while ensuring structural consistency and fidelity. The visualization in Fig. 6 confirms this as well. In the initial stages of denoising, the adapter encourages the integration of ControlNet features by increasing the control weight map. Subsequently, it progressively reduces these control weights to suppress ControlNet constraints, thereby guiding the generation of high-frequency details to enhance the visual quality of results.
@@
It is noteworthy that when applying , not only does the CLIPIQA metric improve, but all no-reference image quality assessment metrics show an increase. We believe that adopting a better perceptual quality as a reward signal could further enhance the results, and this lies beyond the scope of our current research.
@@
Timestep-Aware training Strategy. We conducted further experiments to explore the impact of using different reward functions at various time steps. Specifically, we first added both and across all time steps (0-1000 steps) in the training process, denoted as ”all steps”. As shown in Tab. 5 and Fig. 7 (a), when image rewards are introduced at all time steps, the model becomes unstable during training and fails to generate the correct HR image. Consequently, all reference and non-reference metrics significantly decrease. Similarly, when we applied the inverse training strategy that introduces in the early denoising stages (0-800 steps) and in the later stages (0-200 steps), the model also suffered from instability during training, resulting in lower quality high-resolution images. These experimental results demonstrate that our training strategy is crucial for achieving optimal outcomes.
| Exp | ||||||
| w/o timestep | 20.66 | 0.4969 | 0.3965 | 0.5856 | 67.38 | 0.7668 |
| Transformer | 21.09 | 0.5235 | 0.3619 | 0.5790 | 67.64 | 0.7583 |
| Ours | 20.92 | 0.5174 | 0.3762 | 0.6007 | 68.14 | 0.7681 |
| exps | ||||||
| ControlNet | 19.12 | 0.4344 | 0.4632 | 0.6876 | 74.90 | 0.9345 |
| Adapter | 21.42 | 0.5342 | 0.3561 | 0.5628 | 66.75 | 0.7501 |
| Ours | 20.92 | 0.5174 | 0.3762 | 0.6007 | 68.14 | 0.7681 |
| ✗ | ✗ | 20.88 | 0.5117 | 0.3639 | 0.5781 | 67.63 | 0.7566 |
| ✗ | ✓ | 19.98 | 0.4570 | 0.4110 | 0.6265 | 69.73 | 0.7687 |
| ✓ | ✗ | 21.02 | 0.5224 | 0.3650 | 0.5586 | 67.12 | 0.7366 |
| ✓ | ✓ | 20.92 | 0.5174 | 0.3762 | 0.6007 | 68.14 | 0.7681 |
| exps | ||||||
| all steps | 17.97 | 0.4079 | 0.6299 | 0.5302 | 65.12 | 0.6790 |
| inverse steps | 18.80 | 0.4397 | 0.5830 | 0.5863 | 67.48 | 0.7508 |
| Ours | 20.92 | 0.5174 | 0.3762 | 0.6007 | 68.14 | 0.7681 |
5 Conclusion
@@
In this paper, we proposed a timestep-aware image super-resolution method that introduces a timestep-aware adapter to dynamically integrate ControlNet and diffusion features. In addition, we designed a timestep-aware training strategy to train each module separately based on the generative pattern of the denoising process. Extensive experiments on benchmark datasets demonstrate the effectiveness of our method compared with the state-of-the-art methods.

### v2

@@ 4. Experiments
| RealLR200 | 0.3677 | 0.2968 | 0.2152 | 0.5048 | 0.4324 | 0.4170 | 0.4511 | 0.4673 | 0.6150 | |
| 62.94 | 52.19 | 35.81 | 69.50 | 66.73 | 62.17 | 66.02 | 64.86 | 71.05 | ||
| 0.5392 | 0.3674 | 0.3682 | 0.6811 | 0.6575 | 0.6392 | 0.6760 | 0.6156 | 0.7660 | ||
4.1. Dataset and Evaluation Metric
@@
Datasets. Following (Wu et al., 2024; ao Yang et al., [n. d.]; Lin et al., 2024), we train our method on DIV2K (Agustsson and Timofte, 2017), DIV8K (Gu et al., 2019), Flickr2K (Timofte et al., 2017), OST (Wang et al., 2018a), and the first 10,000 face images from FFHQ (Karras et al., 2019), and use the degradation pipeline of RealESR-GAN (Wang et al., 2021) to generate HR-LR image pairs for training. We evaluate the performance of our method on both synthetic and real-world datasets. The synthetic dataset is generated from the DIV2K validation set, where we use the same degradation pipeline in the training process to randomly crop 3K image patches to as LR images. For real-world test datasets, we evaluate the DrealSR (Wei et al., 2020), RealSR (Wei et al., 2020), and RealLR200 (Wu et al., 2024)datasets. In both DrealSR and RealSR datasets, each image is center-cropped to obtain LR images of resolution. The resolution of each HR image in both the training and test sets is . The RealLR200 dataset is built by SeeSR, which comprises 200 LR images from recent works and a series of real-world scenes.
@@
Metrics. We perform a comprehensive and effective quantitative evaluation of ISR methods using a series of widely used reference and non-reference metrics. Among the reference-based metrics, PSNR and SSIM (Wang et al., 2004) (calculated on the Y channel in the YCbCr space) are fidelity metrics, while LPIPS (Zhang et al., 2018) are quality assessment metrics. MANIQA (Yang et al., 2022), MUSIQ (Yang et al., 2022), and CLIPIQA (Wang et al., 2023) are non-reference image quality assessment (IQA) metrics.
4.2. Implementation Details
@@
We use the pre-trained SD-v2.1 (Rombach et al., 2022b) model as the base SD model, with the ControlNet module being initialized using the base SD model parameters. In the first stage of training, we fine-tune the ControlNet module for 20K iterations, and in the second stage, we fine-tune both the ControlNet and Adapter for 100K iterations with the same training dataset. During the two-stage training, we use the AdamW (Kinga et al., 2015) optimizer with a weight decay of 1e-2, a batch size of 32, and a learning rate of 1e-5. All experiments are conducted at a resolution of 512×512 on 8 NVIDIA A100 GPUs. During inference, we employ a classifier-free guidance strategy, generating higher-quality images through given negative prompts without additional training. The classifier-free guidance scale is set to 4.5, and we use a spaced DDPM sampling schedule (Nichol and Dhariwal, 2021) with 20 time steps.
4.3. Comparison with State-of-the-art Methods
@@
To validate the effectiveness of our method, we compared it with several state-of-the-art GAN-based and diffusion-based ISR methods, namely RealESRGAN (Wang et al., 2021), BSRGAN (Zhang et al., 2021), SwinIR (Liang et al., 2021), SeeSR (Wu et al., 2024), PASD (ao Yang et al., [n. d.]), ResShift (Yue et al., 2024), DiffBIR (Lin et al., 2024), and SUPIR (Yu et al., 2024).
@@
Quantitative Comparisons. Tab. 1 presents the quantitative comparison on both synthetic and real-world test datasets. As observed, our method significantly outperforms other state-of-the-art (SOTA) methods on non-reference metrics such as MANIQA, MUSIQ, and CLIPIQA across all datasets, generating higher-quality HR images. For example, on the DRealSR dataset, TASR outperforms the second-best method, SeeSR by 7.8%, 0.4%, and 5.0% on MANIQA, MUSIQ, and CLIPIQA metrics, respectively. On the RealLR200 dataset, compared to the second-best method, TASR achieves improvements of 21.83% and 12.46% on the MANIQA and CLIPIQA metrics, respectively. Furthermore, GAN-based methods surpass diffusion-based methods on reference metrics such as PSNR and SSIM. We attribute this phenomenon to the fact that diffusion-based methods leverage powerful generative priors to generate details that are more perceptually realistic to humans, but this comes at the expense of fidelity to the LR images. As noted in previous works (Wu et al., 2024; ao Yang et al., [n. d.]), this fundamental trade-off results in high perception quality but low scores on reference metrics. Fig. 4 demonstrates our method can generate high-quality images aligned with human perception while lagging in reference metrics in some scenes.
@@ 4.4. Ablation Study
Model Architecture. We employ various architectures to validate the effectiveness of our proposed model structure for the adapter. Firstly, we removed all time-adaptive normalization layers from the adapter and directly utilized the U-Net features and the skip connection features to predict the control weight map . This modification is denoted as ‘w/o timestep’. As shown in Tab. 2, by incorporating timesteps into the adapter, our method achieves better performance in all the metrics compared to the variant without timesteps. In addition, we replace our adapter structure with a transformer-based architecture similar to ELLA (Hu et al., 2024), denoted as ‘Transformer’. The variant with the Transformer block has declined in the non-reference metrics.
| w/o timestep | 0.3965 | 0.5856 | 67.38 | 0.7668 |
| Transformer | 0.3619 | 0.5790 | 67.64 | 0.7583 |
| ControlNet | 0.4632 | 0.6876 | 74.90 | 0.9345 |
| Adapter | 0.3561 | 0.5628 | 66.75 | 0.7501 |
| Ours | 0.3762 | 0.6007 | 68.14 | 0.7681 |
@@
Training for Different Module. Firstly, we remove the timestep-aware adapter and optimize only the parameters of ControlNet during the training process, denoted as ‘ControlNet’. As shown in Tab. 2 and Fig. 6, when the limitation of the adapter on the optimization space is removed, the method is prone to reward hacking during training. Although there is a significant improvement in perceptual metrics such as CLIPIQA and MUSIQ, the generated images tend to align with the preferences of these metrics rather than human perception. This can result in images that are not sufficiently realistic and lack fidelity, such as the strange feathers and the human eye in the sky background in Fig. 6 (a). In addition, we employ GPT-4o for comprehensive image quality assessment to determine which model produces images with enhanced naturalness and realism, while utilizing Gram matrices to quantify style discrepancies between generated results and ground truth. Compared to the method without the adapter, our method demonstrates superior performance with GPT-4o that prefers our outputs in 84% of test cases. Furthermore, we achieve a 40.4% reduction in Gram matrix loss from 0.3051 to 0.1817 relative to GT images.
Similarly, we add the proposed adapter but optimize only the parameters of the adapter during the training process, denoted as ‘Adapter’. Our method shows a better trade-off between the reference and non-reference metrics. When only the adapter is optimized, the parameter space available for optimization is constrained, resulting in a 6.3% and 2.0% decrease in the MANIQA and MUSIQ metrics, respectively. As can be seen from Fig. 6, compared to our method, the model obtained by optimizing only the adapter struggles to generate more refined feathers.
@@
Loss Function. We begin by evaluating the impact of the chosen loss functions on the results. As shown in Tab. 3, when only is added as the loss function (row 2), the no-reference image quality assessment metrics MANIQA improves by 8.3% compared to using only the denoising loss (row 1). However, reference metrics such as PSNR and SSIM decrease by 4.3% and 10.6% respectively, indicating lower fidelity in the generated HR images. Similarly, when only is added as the loss function (row 3), the reference metrics improve while the non-reference metrics correspondingly decline. In contrast, our proposed training strategy, applying and , allows us to enhance image perceptual quality while ensuring structural consistency and fidelity. The visualization in Fig. 7 confirms this as well. In the initial stages of denoising, the adapter encourages the integration of ControlNet features by increasing the control weight map. Subsequently, it progressively reduces these control weights to suppress ControlNet constraints, thereby guiding the generation of high-frequency details to enhance the visual quality.
| 0.3639 | 0.5781 | 67.63 | 0.7566 | ||
| 0.4110 | 0.6265 | 69.73 | 0.7687 | ||
| 0.3650 | 0.5586 | 67.12 | 0.7366 | ||
| 0.3762 | 0.6007 | 68.14 | 0.7681 |
@@
It is noteworthy that when applying , not only does the CLIPIQA metric improve, but all no-reference image quality assessment metrics show an increase. During training, both MANIQA and MUSIQ metrics can be employed as image rewards. As shown in Tab. 4, where ‘MANIQA’ denotes that the reward function in Equation (5) corresponds to the MANIQA metrics. Compared to MUSIQ and MANIQA, CLIPIQA results in a more stable training process and better trade-off. Our implementation of CLIPIQA substantiates incorporating the quality metrics, as the image rewards can improve perceptual quality. We believe that adopting a better perceptual quality as a reward signal could further enhance the results, but this lies beyond the scope of our research.
| all steps | 0.6299 | 0.5302 | 65.12 | 0.6790 |
| inverse steps | 0.5830 | 0.5863 | 67.48 | 0.7508 |
| MANIQA | 0.4742 | 0.6105 | 67.93 | 0.7572 |
| MUSIQ | 0.4090 | 0.6241 | 69.23 | 0.7651 |
| CLIPIQA | 0.3762 | 0.6007 | 68.14 | 0.7681 |
@@
Timestep-Aware training Strategy. We conducted further experiments to explore the impact of using different reward functions at various time steps. Specifically, we first added both and across all time steps (0-1000 steps) in the training process, denoted as ”all steps”. As shown in Tab. 4 and Fig. 8 (a), when image rewards are introduced at all time steps, the model becomes unstable during training and fails to generate the correct HR image. Consequently, all reference and non-reference metrics significantly decrease. Similarly, when we applied the inverse training strategy that introduces in the early denoising stages (0-800 steps) and in the later stages (0-200 steps), denoted as ”inverse steps”, the model also suffered from instability during training, resulting in lower quality high-resolution images. These experimental results demonstrate that our training strategy is crucial for achieving optimal outcomes.
5. Conclusion
@@
In this paper, we proposed a timestep-aware image super-resolution method that introduces a timestep-aware adapter to integrate ControlNet and diffusion features dynamically. In addition, we designed a timestep-aware training strategy to train each module separately based on the generative pattern of the denoising process. Extensive experiments on benchmark datasets demonstrate the effectiveness of our method compared with the state-of-the-art methods.
Acknowledgments
This research was funded through National Key Research and Development Program of China (Project No. 2022YFB36066), in part by the Shenzhen Science and Technology Project under Grant (KJZD20240903103210014, JCYJ20220818101001004).